OpenStudy (anonymous):

A health food store sells oatmeal for $3.60 per pound and bran flakes for $4.80 per pound. How many pounds of each should be used to get a mixture of 30 pounds that sells for $4.00 a pound?

OpenStudy (anonymous):

x+y=30 3.6x+4.8y=30*4 =>x=20 y=10